OpenStudy (toxicsugar22):

Dichloro is a pesticide that was used in the middle decades of the 20th century to control maleria. Afterr 1945, it was also widely used on crops in the U, and as much as one ton be sprayed on a single cotton field. Hoowever after the toxic, effects of Dichloro on the evvironment began to appear, the chemical was banned in 1972. A common estimate for the half-life in the soil is 15 years. Write a decay law for Dichloro in the soil and use A for the initial concentration of Dicloro, and use t for the input variable

OpenStudy (anonymous):

N = No*e^(-ht) where h is the half life, so: N = A * e^(-15t)

OpenStudy (toxicsugar22):

that the answer

OpenStudy (anonymous):

No, I made an error

OpenStudy (anonymous):

ln(.5) = -15h so ln(2) = 15h and h = ln(2)/15 so...

OpenStudy (anonymous):

N=A*e^(-ln(2)/15*t) note that t is not in the denominator

OpenStudy (toxicsugar22):

the answere is not clear

OpenStudy (anonymous):

N=A*e^(-t*ln(2)/15)